# Mining the Minoria: Unknown, Under-represented, and Under-performing Minority Groups

### v1

@@ Abstract.
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
To address this problem, we introduce the problem of mining the potential groups that are under-represented and under-performing, when the grouping attributes are unknown. We refer to this problem as Minoria mining. To the best of our knowledge, this paper is the first to introduce and study the Minoria mining problem. In simple terms, the main question we want to answer in this paper is, “are there unknown groups that are (a) under-represented, for which (b) the model under-performs?”
@@ 1.1. Technical Contributions
Consider projecting data on a combination [^1] We interchangeably use the term “linear direction” or simply “direction” for “linear combination”. of features, such that (i) the projected data is highly skewed (hence its tail indicating a potentially under-represented group) and (ii) the prediction loss at the tail is high. Our mining task is to discover such cases and return the under-performing samples (e.g., the misclassified samples) in the tail to the human expert for further investigation to verify if those represent a socially-valid group.
Figure 1. Toy example: Minoria mining using high-skew projections. The orange area shows the tail of the high-skew direction.
To better explain our approach, let us consider a toy example of a basketball dataset with several observation features $X$, including height and salary, and one target attribute ($y$) performance-score, where the goal is to predict $y$ using $X$. Let the (height, salary) values be as shown in Fig. 1. Our system discovers the red vector as a combination of height and salary, where the projection of values on it are highly skewed. That is, after orthogonally projecting each point on the red vector, the distribution of the projected values is highly skewed (in this example most values are large). We consider the tail of projected data (the left region highlighted in orange) as a potentially under-represented group. Next, we observe that the prediction performance is significantly less for this potential group. We then report the misclassified samples on the tail as a potential Minoria to the user for further investigation. In this example, the discovered group represents the female athletes that are not well-represented in the dataset.
@@ 2.1. Problem Definition
Addressing Problem 1 is challenging as it requires identifying the median of each possible projection while there are infinite possible projections defined using different vectors $f$. In order to tackle this challenge, we provide a geometric dual transformation of $\mathcal{D}$ in the next section that helps us design our efficient algorithm in § 4.
@@ 3.2. Median Region
Now consider the origin-anchored ray $\vec{r_{f}}$ in the dual space and the intersections of the dual hyperplanes with it. Based on Lemma 1, the $k$-th closest intersection to the origin specifies the tuple with the $k$-th largest value in $\mathcal{D}_{f}$. We define the universe of all such intersection points for all vectors $f$ as the $k$-th level of arrangement. We refer to the $k$-th level of the arrangement of $\mathcal{H}$ as $\mathcal{A}_{k}$. In $\mathbb{R}^{2}$, we show $\mathcal{A}_{k}$ as the ordered list of the vertices (the intersection points) of the $k$-th level, in counter-clock-wise (from $x$-axis to $y$) or clock-wise.
@@
Using these observations, one can write a linear program for every median region to find the maximum skew within a median region. Instead, we introduce Theorem 3 as a key observation that enables a more efficient algorithm in Section 4.
Let $\mu(\mathcal{D})$ be the vector in $\mathbb{R}^{d}$ that contains the mean of $\mathcal{D}$, i.e., $\mu(\mathcal{D})=\frac{1}{n}\sum_{i=1}^{n}t_{i}$. We define $q_{i}=t_{i}-\mu(\mathcal{D})$. In other words, $q_{i}$’s are the mean-normalized tuples of $\mathcal{D}$. Let $Q$ be a $d\times n$ matrix where each column of $Q$ is $q_{i}^{\top}$. For a unit vector $f$, let $t_{m_{f}}$ be the median of $\mathcal{D}$ after projection on $f$. In other words, $t_{m_{f}}^{\top}f$ is the median of $\mathcal{D}_{f}$. The following theorem helps us to find the high skew direction vector $f^{*}$ efficiently:
@@
Proof.
First, let us calculate the absolute value of $skew$ for an arbitrary $f$:
$skew(f)$ $=\Big{\lvert}\frac{mean(\mathcal{D}_{f})-median(\mathcal{D}_{f})}{sd(\mathcal{D}_{f})}\Big{\rvert}$
$=\Big{\lvert}\frac{\mu(\mathcal{D})^{T}f-t_{m_{f}}^{T}f}{sd(\mathcal{D}_{f})}\Big{\rvert}=\Big{\lvert}\frac{(\mu(\mathcal{D})-t_{m_{f}})^{T}f}{sd(\mathcal{D}_{f})}\Big{\rvert}$
The $sd(\mathcal{D}_{f})$ is as follows:
$sd(\mathcal{D}_{f})$ $=\sqrt{\frac{\sum^{n}_{i=1}(mean(\mathcal{D}_{f})-t_{i}^{T}f)^{2}}{n}}=\sqrt{\frac{\sum^{n}_{i=1}((\mu(\mathcal{D})-t_{i})^{T}f)^{2}}{n}}$
Now, replacing $\mu(\mathcal{D})-t_{i}$ with $q_{i}$s, we have:
$skew(f)$ $=\Big{\lvert}\frac{q_{m_{f}}^{T}f}{\sqrt{\frac{\sum^{n}_{i=1}(q_{i}^{T}f)^{2}}{n}}}\Big{\rvert}$
Our objective is to find the $f^{*}=\operatorname*{arg\,max}$ of $skew$:
$f^{*}$ $=\operatorname*{arg\,max}_{f}skew(f)=\operatorname*{arg\,max}_{f}skew(f)^{2}$
$=\operatorname*{arg\,max}_{f}\frac{(q_{m_{f}}^{T}f)^{2}}{\sum^{n}_{i=1}(q_{i}^{T}f)^{2}}=\operatorname*{arg\,max}_{f}\frac{(q_{m_{f}}^{T}f)^{2}}{f^{T}QQ^{T}f}$
Where $Q$ is the matrix with $q_{i}$s as the columns. The vector $Q^{T}f$ will have $q_{i}^{T}f$ as its elements. The dot product of this vector and its transpose will give $\sum_{i=1}^{n}(q_{i}^{T}f)^{2}$. Let us define $g(f)=\frac{(q_{m_{f}}^{T}f)^{2}}{f^{T}QQ^{T}f}$.
The derivative of the $g(f)$ with respect to $f$ is:
$\frac{\partial g(f)}{\partial f}$ $=\frac{\partial(q_{m_{f}}^{T}f)^{2}}{\partial f}f^{T}QQ^{T}f-\frac{\partial(f^{T}QQ^{T}f)}{\partial f}(q_{m_{f}}^{T}f)^{2}$
$=2q_{m_{f}}^{T}fq_{m_{f}}(f^{T}QQ^{T}f)-2QQ^{T}f(q_{m_{f}}^{T}f)^{2}$
Solving the equation $\frac{\partial g(f^{*})}{\partial f}=0$, we get:
$\frac{\partial g(f^{\star})}{\partial f}=0$ $\Leftrightarrow 2(q_{m_{f}}^{T}f^{\star})(f^{\star T}QQ^{T}f^{\star})q_{m_{f}}=2(q_{m_{f}}^{T}f^{\star})^{2}QQ^{T}f^{\star}$
$\Leftrightarrow(f^{\star T}QQ^{T}f^{\star})q_{m_{f}}=(q_{m_{f}}^{T}f^{\star})QQ^{T}f^{\star}$
$\Leftrightarrow f^{\star}=c(QQ^{T})^{-1}q_{m_{f}}$
We assumed $q_{m_{f}}^{T}f^{\star}\neq 0$ to calculate the $\operatorname*{arg\,max}$ of $skew$. In the last line, $c$ is a scalar value. So the direction of $f^{*}$ would be equal to $(QQ^{T})^{-1}q_{m_{f}}$. The unit vector for this direction would be $\frac{(QQ^{T})^{-1}q_{m_{f}}}{|(QQ^{T})^{-1}q_{m_{f}}|}$.
So far, we have shown that, for each median region, the maximum skew direction $f^{*}$ happens in the single direction $(QQ^{\top})^{-1}q_{m_{f}}$. However, this direction may fall outside this median region. In such cases, the median for this direction will be a different tuple. Hence, the argument about the maximum skew is no longer valid. For such cases, the maximum-skew direction $f^{*}$ is one of the boundary points of the median region. That is because the $skew$ function is monotonic inside each median region. ∎
@@ 4. Minoria Mining in 2D
In the end, we pop the $-\ell$ directions from the heap, evaluate the model on the tail of these directions, and check the error on the tail to see if it represents an under-represented group candidate.
@@ 4.2. The Warm-up Algorithm
The Warm-up algorithm (Algorithm 1) conducts the three aforementioned steps separately. In the first step, it transforms the dataset into the dual hyperplanes and calls the oracle $\mathcal{E}$ to enumerate $\mathcal{A}_{\frac{n}{2}}$. As a result, the first step takes $T_{enum}=O(n\log(nk^{\frac{1}{3}})+nk^{\frac{1}{3}}\log^{1+\alpha}n)=O(n^{\frac{4}{3}}\log^{1+\alpha}n)$ for the deterministic algorithm, or $T_{enum}=O(nk^{\frac{1}{3}}+n\log n)=O(n^{\frac{4}{3}})$ when using the randomized incremental algorithm (note that $k=\frac{n}{2}$ here).
In the second step, for each vertex in $\mathcal{A}_{\frac{n}{2}}$, it calculates the $skew$ value and adds it to a max-heap. Calculating the skew requires calculating the mean, median, and standard deviation of points projected on the direction $f$ and takes $O(n)$. Inserting into a max-heap using a Fibonacci heap takes a constant time. As a result, $T_{skew}=O(n^{2}k^{\frac{1}{3}})=O(n^{\frac{7}{3}})$.
In the third step, the Warm-up algorithm pops the high skew directions from the heap and calculates the model’s loss on the tail of each direction. Popping from a max-heap takes an amortized time of $O(\log n)$. Calculating the model accuracy would take $O(n)$. Hence, assuming that the number of directions evaluated before returning the $-\ell$ is a small constant, $T_{pop}=O(n\,\log n)$.
@@
Algorithm 1 Warm-up Algorithm for 2D
1:procedure 2DWarm-up($\mathcal{D},\theta$)
2:     $\mathcal{H}\leftarrow Dual(\mathcal{D})$
3:     $\mathcal{A}_{\frac{n}{2}}\leftarrow\mathcal{E}(\mathcal{H},\frac{n}{2})$$\triangleright$ Ordered list of vertices in $\frac{n}{2}$th level of arrangement.
4:     $Heap\leftarrow[~{}]$$\triangleright$ Initialize a max heap.
5:     for $v$ in $\mathcal{A}_{\frac{n}{2}}$ do
6:         $f\leftarrow v/||v||$$\triangleright$ Get the direction of point $v$.
7:         $skew\leftarrow GetSkewNaive(\mathcal{D},f)$
8:         $Heap.push(\langle f,skew\rangle)$$\triangleright$ Push $f$ to a max heap based on the skew value.
9:     $Output\leftarrow[~{}]$
10:     while $Output$ size is less that $l$ do
11:         $f\leftarrow Heap.pop()$$\triangleright$ Get a candidate direction.
12:         $T\leftarrow\textit{p-}tail(f)$$\triangleright$ $T$ is a candidate Minoria subset of $\mathcal{D}$.
13:         if $L_{T}(\theta)-L_{\mathcal{D}}(\theta)>\tau$ then $Output.append(f)$
14:     return $Output$$\triangleright$ candidate Minoria directions.
15:procedure GetSkewNaive($\mathcal{D},f$)
16:     return $3(\mu(\mathcal{D}_{f})-\nu(\mathcal{D}_{f}))/\sigma(\mathcal{D}_{f})$
@@ 4.3. The Ray-sweeping Algorithm
The Warm-up algorithm is super-quadratic and not efficient. Therefore, in this section, we propose the Ray-sweeping algorithm that has the same time complexity as of the enumeration of $\mathcal{A}_{\frac{n}{2}}$. At a high level, we maintain some aggregates while sweeping over $\mathcal{A}_{\frac{n}{2}}$ from the x-axis to the y-axis, which enables updating the skew values in constant time.
The key observation is that while moving between two neighboring median regions, the set of dual lines above/below the median changes by only one element. As a result, using the pre-calculated values for the previous vertex, we can calculate the skew value of the current vertex in constant time (except for the very first vertex in the list). Please Refer to § 5.2 for more details about the constant-time skew-value update process. Moreover, when visiting a new vertex in $\mathcal{A}_{\frac{n}{2}}$, the new median can be found in constant time simply because each vertex is the intersection of the medians of the two neighboring median regions.
@@ 5.1. Generalized Ray-sweeping Algorithm
2:     $\mathcal{H}\leftarrow Dual(\mathcal{D})$
@@
7:         $skew=GetSkew(\mathcal{D},f)$$\triangleright$ Use the aggregated values to get skew in constant time.
@@
9:     $Output\leftarrow[~{}]$
10:     while $Output$ size is less that $l$ do
11:         $f\leftarrow Heap.pop()$
12:         $T\leftarrow\textit{p-}tail(f)$
13:         if $L_{T}(\theta)-L_{\mathcal{D}}(\theta)>\tau$ then $Output.append(f)$
14:     return $Output$
@@ 6.1.2. The Candidate-group diversification Approach
6.1.3. Error Bound (Negative Result)
In this section, we analyze the error bounds of the discretization approaches. We provide a negative result that an upper bound on the ratio of the skew of a direction $f^{\prime}$ with a cosine similarity at least $cos(\alpha)$ with the best direction $f^{*}$ depends on the dataset topology.
First, let us find an upper bound on the inner product of two vectors while rotating one of them. Given a fixed vector $\vec{a}=(a_{1},a_{2})$ and a variable vector $\vec{f}$. We rotate $\vec{f}=(m,n)$ by a rotation matrix in 2D space $R_{\alpha}$ ($\alpha$ is the angle):
(9) $\hskip 56.9055ptR_{\alpha}=\begin{bmatrix}\cos{\alpha}&-\sin(\alpha)\\ \sin{\alpha}&\cos{\alpha}\end{bmatrix}$
Now, calculating the ratio $\frac{\vec{a}^{\top}(R_{\alpha}\vec{f})}{\vec{a}^{\top}\vec{f}}$, we get:
$\frac{\vec{a}^{\top}(R_{\alpha}\vec{f})}{\vec{a}^{\top}\vec{f}}$ $=\frac{a_{1}(m\cos(\alpha)-n\sin(\alpha))+a_{2}(m\sin(\alpha)+n\cos(\alpha))}{a_{1}m+a_{2}n}$
$=\frac{(a_{1}m+a_{2}n)\cos(\alpha)+(-a_{1}n+a_{2}m)\sin(\alpha)}{a_{1}m+a_{2}n}$
$=\cos(\alpha)+\sin(\alpha)\frac{a_{2}m-a_{1}n}{a_{1}m+a_{2}n}=\cos(\alpha)+\sin(\alpha)\frac{\vec{a_{\perp}}^{\top}\vec{f}}{\vec{a}^{\top}\vec{f}}$
$=\cos(\alpha)+\sin(\alpha)\tan(\beta)$
Where $\vec{a_{\perp}}$ is the rotation of $\vec{a}$ as $90$ degrees counter clockwise and $\beta$ is the angle between $\vec{a}$ and $\vec{f}$.
To find an upper bound on the skew of a vector when it is rotated with a small $\alpha$ angel, we consider the following:
(10) $Skew(f)=\frac{3(mean(\mathcal{D}_{f})-median(\mathcal{D}_{f}))}{sd(\mathcal{D}_{f})}$ $=\frac{3q_{m_{f}}^{\top}f}{\sqrt{\sum_{i}(q_{i}^{\top}f)^{2}}}$
Where $q_{i}=mean(P)-p_{i}$ and $m_{f}$ is the index of the median point. Now we rotate $f$ by the rotation matrix $R_{\alpha}$. Assume the median does not change in this rotation.
(11) $\frac{Skew(R_{\alpha}f)}{Skew(f)}$ $=\frac{q_{m_{f}}^{\top}(R_{\alpha}f)}{q_{m_{f}}^{\top}f}\sqrt{\frac{\sum_{i}(q_{i}^{\top}f)^{2}}{\sum_{i}(q_{i}^{\top}(R_{\alpha}f))^{2}}}$
Now using the bound for Inner Product ($\beta_{i}$ is the angle between $f$ and $q_{i}$.):
(12) $\frac{q_{m_{f}}^{\top}(R_{\alpha}f)}{q_{m_{f}}^{\top}f}\sqrt{\frac{\sum_{i}(q_{i}^{\top}f)^{2}}{\sum_{i}(q_{i}^{\top}(R_{\alpha}f))^{2}}}$
$=(\cos(\alpha)+\sin(\alpha)\tan(\beta_{m}))\sqrt{\frac{\sum_{i}(q_{i}^{\top}f)^{2}}{\sum_{i}(q_{i}^{\top}R_{\alpha}f)^{2}}}$
$\leq\frac{\cos(\alpha)+\sin(\alpha)\tan(\beta_{m})}{\cos(\alpha)+\sin(\alpha)\tan(\beta_{min})}=\frac{1+\tan(\alpha)\tan(\beta_{m_{f}})}{1+\tan(\alpha)\tan(\beta_{min})}$
$\leq 1+\tan(\alpha)\tan(\beta_{m_{f}})$
Where $\beta_{min}=\min_{i}{\beta_{i}}$. This shows that the upper bound on the skew ratio depends on the angle between the direction $f$ and the median point $q_{m_{f}}$ in the normalized dataset. Therefore, we cannot find an upper bound independent of the dataset topology.
As a result, when we consider the discretized space of all possible directions which are at most $\epsilon$ angle apart, the error bound on the $skew$ value depends on the topological characteristics of the dataset. If the median point is far from the chosen function $f$, based on the above calculations, moving the $f$ slightly may significantly change the $skew$.
Even though we cannot guarantee the change in the skew value of resulting directions of the discretization approaches, we guarantee that for each direction $f$ in the $-\ell$ set of local optima high-skew directions (Refer to Problem Definition  1), we can find a direction $f^{\prime}$ such that $cos(f^{\prime},f)\geq\epsilon$ where $\epsilon$ is the hyper-parameter used in the grid partitioning approach.

### v2

@@ Abstract.
PVLDB, 18(5): 1453 - 1465, 2025.
doi:10.14778/3718057.3718072 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 5 ISSN 2150-8097.
doi:10.14778/3718057.3718072
@@ 1. Introduction
To address this problem, we introduce the problem of mining the potential groups that are under-represented and under-performing, when the grouping attributes are unknown. We refer to this problem as Minoria mining. To the best of our knowledge, this paper is the first to introduce and study the Minoria mining problem. In simple terms, the main question we want to answer in this paper is, “are there regions in data-space, representing potential groups that are (a) under-represented, for which (b) the model under-performs?” For instance, aligned with Example 1, our experiments on a Crimes dataset (§ 7.1.1) identified North-side Chicago as under-represented and under-performing. This information helps the model developers and data owners such as Chicago Open Data Portal to proactively identify a potential issue, decide if it is socially concerning, and limit the dataset’s scope-of-use accordingly.
@@ 1.1. Technical Contributions
Consider projecting data on a combination [^1] We interchangeably use the term “linear direction” or simply “direction” for “linear combination”. of features, such that (i) the projected data is highly skewed (hence its tail indicating an under-represented group candidate) and (ii) the prediction loss at the tail is high. Our mining task is to discover such cases and return the under-performing samples (e.g., the misclassified samples) in the tail to the human expert for further investigation to verify if those represent a socially-valid group.
Figure 1. Toy example: Minoria mining using high-skew projections.
To better explain our approach, let us consider a toy example of a basketball dataset with several observation features $X$, including height and salary, and one target attribute ($y$) performance-score, where the goal is to predict $y$ using $X$. Let the (height, salary) values be as shown in Fig. 1. Our system discovers the black vector as a combination of height and salary, where the projection of values on it are highly skewed. That is, after orthogonally projecting each point on the black vector, the distribution of the projected values is highly skewed. In this example, the projections are specified with dashed line, while the point in the head and tail of the distribution colored in red and blue, respectively. The points in the tail are considered as a potential groups that are under-represented. Next, we observe that the prediction performance is significantly less for this potential group. We then report the misclassified samples on the tail as a potential Minoria to the user for further investigation. In this example, the discovered group represents (mostly) the female athletes that are not well-represented in the dataset.
@@ 2.1. Problem Definition
Addressing Problem 1 is challenging as it requires identifying the median of each possible projection while there are infinite possible projections defined using different vectors $f$. To tackle this challenge, we provide a geometric dual transformation of $\mathcal{D}$ in the next section that helps us design our efficient algorithm in § 4.
@@ 3.2. Median Region
Now consider the origin-anchored ray $\vec{r_{f}}$ in the dual space and the intersections of the dual hyperplanes with it. Based on Lemma 1, the $k$-th closest intersection to the origin specifies the tuple with the $k$-th largest value in $\mathcal{D}_{f}$. We define the universe of all such intersection points for all vectors $f$ as the $k$-th level of arrangement, referred as $\mathcal{A}_{k}$. In $\mathbb{R}^{2}$, we show $\mathcal{A}_{k}$ as the ordered list of the vertices (the intersection points) of the $k$-th level, in counter-clock-wise (from $x$-axis to $y$) or clock-wise.
@@
Using these observations, one can write a linear program for every median region to find the maximum skew within a median region. Instead, we introduce Theorem 3 as a key observation that enables a more efficient algorithm in Section 4. Let $\mu(\mathcal{D})$ be the vector in $\mathbb{R}^{d}$ that contains the mean of $\mathcal{D}$, i.e., $\mu(\mathcal{D})=\frac{1}{n}\sum_{i=1}^{n}t_{i}$. We define $q_{i}=t_{i}-\mu(\mathcal{D})$. In other words, $q_{i}$’s are the mean-normalized tuples of $\mathcal{D}$. Let $Q$ be a $d\times n$ matrix where each column of $Q$ is $q_{i}^{\top}$. For a unit vector $f$, let $t_{m_{f}}$ be the median of $\mathcal{D}$ after projection on $f$. In other words, $t_{m_{f}}^{\top}f$ is the median of $\mathcal{D}_{f}$. Theorem 3 helps us to find the high skew direction vector $f^{*}$ efficiently [^4] Due to the space limitations, the proof is provided in the technical report [16].:
@@ 4. Minoria Mining in 2D
In the end, we pop the $-\ell$ directions from the heap, evaluate the model on the tail of these directions, and check the error on the tail to see if it represents an under-represented group candidate. A visualization of this approach is provided in Figure 3.
@@ 4.2. The Warm-up Algorithm
So far, we outlined the key steps in our algorithmic framework. Following these steps, the Warm-up algorithm [^5] Please find the pseudo-code of the Warm-up algorithm in the technical report [16]. first enumerates $\mathcal{A}_{\frac{n}{2}}$. It then computes the median by evaluating all projections along a specified direction. Finally, it identifies and retrieves the top skewed elements.
Complexity Analysis
The Warm-up algorithm conducts the three aforementioned steps separately. In the first step, it transforms the dataset into the dual hyperplanes and calls the oracle $\mathcal{E}$ to enumerate $\mathcal{A}_{\frac{n}{2}}$. As a result, the first step takes $T_{enum}=O(n\log(nk^{\frac{1}{3}})+nk^{\frac{1}{3}}\log^{1+\alpha}n)=O(n^{\frac{4}{3}}\log^{1+\alpha}n)$ for the deterministic algorithm, or $T_{enum}=O(nk^{\frac{1}{3}}+n\log n)=O(n^{\frac{4}{3}})$ when using the randomized incremental algorithm (note that $k=\frac{n}{2}$ here). In the second step, for each vertex in $\mathcal{A}_{\frac{n}{2}}$, it calculates the $skew$ value and adds it to a max-heap. Calculating the skew requires calculating the mean, median, and standard deviation of points projected on the direction $f$ and takes $O(n)$. Inserting into a max-heap using a Fibonacci heap takes a constant time. As a result, $T_{skew}=O(n^{2}k^{\frac{1}{3}})=O(n^{\frac{7}{3}})$. In the third step, the Warm-up algorithm pops the high skew directions from the heap and calculates the model’s loss on the tail of each direction. Popping from a max-heap takes an amortized time of $O(\log n)$. Calculating the model accuracy would take $O(n)$. Hence, assuming that the number of directions evaluated before returning the $-\ell$ is a small constant, $T_{pop}=O(n\,\log n)$.
@@ 4.3. The Ray-sweeping Algorithm
The Warm-up algorithm is super-quadratic and not efficient. Therefore, in this section, we propose the Ray-sweeping algorithm that has the same time complexity as of the enumeration of $\mathcal{A}_{\frac{n}{2}}$. Similar to the Warm-up algorithm, the Ray-sweeping algorithm also follows the three steps of (1) enumerating $\mathcal{A}_{\frac{n}{2}}$, followed by (2) computing the medians and (3) identifying the high-skew projections. However, unlike Warm-up, it leverages our earlier observation to compute medians more efficiently by sweeping a ray counterclockwise from the x-axis to the y-axis. At a high level, it maintains some aggregates while sweeping over $\mathcal{A}_{\frac{n}{2}}$, which enables updating the skew values in constant time.
The key observation is that while moving between two neighboring median regions, the set of dual lines above/below the median changes by only one element. As a result, using the pre-calculated values for the previous vertex, we can calculate the skew value of the current vertex in constant time (except for the very first vertex in the list). Please Refer to § 5.2 for more details about the constant-time skew-value update process. Moreover, when visiting a new vertex in $\mathcal{A}_{\frac{n}{2}}$, the new median can be found in constant time simply because each vertex is the intersection of the medians of the two neighboring median regions. The pseudo-code of this approach is provided in Algorithm 1. A visualization of this algorithm on a toy example is provided in Figure 3.
Complexity Analysis
@@
(a) An example of $5$ points and their dual lines in the same plane.
(b) The $k$-th level of arrangement for $k=\frac{n}{2}$, representing the median points.
(c) The median regions and their boundaries for a ray starting from the origin.
(d) Sweeping a ray ($\vec{r_{f}}$) from $x$-axis to $y$-axis counterclockwise.
Figure 3. A toy example for the Ray-Sweeping algorithm. (a) and (b): Starting from $5$ points in the plane, the algorithm finds the $\frac{n}{2}$-the level of arrangements for the dual lines. (c): Each median region is a line segment in this arrangement. The line segment represents the median point for any direction in this region. (d): The algorithm sweeps a ray from the $x$-axis to the $y$-axis for each region, calculates the skew value in this region efficiently, and pushes this value into a heap. In the end, the best directions, in terms of their skew, are extracted from the heap.
@@ 5.1. Generalized Ray-sweeping Algorithm
2:    $\mathcal{H}\leftarrow Dual(\mathcal{D})$; $Output\leftarrow[~{}]$
@@
7:         $skew=GetSkew(\mathcal{D},f)$$\triangleright$ Use the aggregated values to compute the skew in constant time.
@@
9:    while $Output$ size is less that $l$ do
10:         $f\leftarrow Heap.pop()$
11:         $T\leftarrow\textit{p-}tail(f)$
12:         if $L_{T}(\theta)-L_{\mathcal{D}}(\theta)>\tau$ then $Output.append(f)$
13:    return $Output$
@@ 6.1.2. The Candidate-group diversification Approach
In our technical report [16], we analyze the error bounds of the discretization approaches and provide the following negative result: let $f^{\prime}$ be a direction with a cosine similarity of (at least) $cos(\alpha)$ with the best direction $f^{*}$. An upper-bound on skew ratio $\frac{f^{\prime}}{f^{*}}$ depends on the dataset topology. As a result, when discretizing the directions with $\epsilon$-angle spacing, the error bound on the $skew$ value depends on the topological characteristics of the dataset. Still, we guarantee that for each direction $f$ in the $-\ell$ high-skew local optima (Refer to Problem Definition  1), we evaluate a direction $f^{\prime}$ such that $cos(f^{\prime},f)\geq\epsilon$.